OpenStudy (anonymous):

1. On a summer day, you take a road trip through Death Valley, California, in an antique car. You start out at a temperature of 21°C, but the temperature in Death Valley will reach a peak of 51°C. The tires on your car hold 15.6 L of nitrogen gas at a starting pressure of 249 kPa. The tires will burst when the internal pressure (Pb) reaches 269 kPa. Answer the following questions and show your work.

OpenStudy (anonymous):

• How many moles of nitrogen gas are in each tire? • What will the tire pressure be at peak temperature in Death Valley? • Will the tires burst in Death Valley? Explain. • If you must let nitrogen gas out of the tire before you go, to what pressure must you reduce the tires before you start your trip? (Assume no significant change in tire volume.)

OpenStudy (anonymous):

aaronq

OpenStudy (anonymous):

aaronq can you help

OpenStudy (aaronq):

"How many moles of nitrogen gas are in each tire?" use the ideal gas law PV=nRT "What will the tire pressure be at peak temperature in Death Valley?" use \(\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}\) "Will the tires burst in Death Valley? Explain." does the pressure go past 269 kPa?

OpenStudy (anonymous):

see i tried that but its confusing to me

OpenStudy (anonymous):

im not sure how that is done i never really picked it up well in my class and my teacher cant seem to explain it that well for me to understand :/ ive tried to do it that way but i get lost and end up making it a big mess

OpenStudy (aaronq):

to find the number of moles, we first determine what each variable is, use the first set of data (not at the canyon). first, you need temp in kelvin: T=273.15+ 21°C = 294.15 K V= 15.6 L P= 249 kPa R= (these you look up to match the units) http://en.wikipedia.org/wiki/Gas_constant =8.314 L*kPa/K*mol so rearrange the formula for moles, n: \(PV=nRT \rightarrow n=\dfrac{PV}{RT}\) \(n=\dfrac{(249 kPa)*(15.6 L)}{(8.314 L*kPa/K*mol)*(294.15~K)}=1.5883 ~moles~ of ~N_2\)

OpenStudy (aaronq):

think you can try the second part?

OpenStudy (anonymous):

so then it is 8.3140 kPa L/(K mol) (kilopascal liters per kelvin mole)?

OpenStudy (aaronq):

yeah, notice i chose it because it matches the pressure (kPa) and volume (L).

OpenStudy (anonymous):

okay lets see if i can do the second one :)

OpenStudy (aaronq):

sounds good. first determine which variable is which.

OpenStudy (anonymous):

Peak temperature in Death Valley is 51 C. However, you can only use proportions with Kelvin temperature so you need to convert Celsius to Kelvin by adding 273. You have to set up another proportion. The volume and temperature of a confined sample of gas are directly proportional. P1/T1 = P2/T2. The temperatures MUST be in Kelvin. So plug in your numbers. 249 kPa/ (21 + 273) K = P2/ (51 + 273) Solve that. If I did the math right, the pressure at the peak temperature would be 274 kPa. Since the peak temperature in Death Valley will cause the pressure in the tires to expand to 274 kPa which exceeds the maximum pressure value, the tires will burst. 269 kPa/ 324 = p2/ 294 p2= 245 kPa. would that be correct?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

awesome! okay then now im confused on the the third point how would i say that or i mean how would i explain it

OpenStudy (anonymous):

just say Since the peak temperature in Death Valley will cause the pressure in the tires to expand to 274 kPa which exceeds the maximum pressure value, the tires will burst.

OpenStudy (anonymous):

for the last point I dont understand whats so ever i feel like the question doesnt give me enough info to solve

OpenStudy (aaronq):

I'm kinda confused with what you posted earlier, the answer is \(P_2\)=274.40 kPa so this "269 kPa/ 324 = p2/ 294 p2= 245 kPa. would that be correct?" is not correct.

OpenStudy (anonymous):

:c se i told you im horrible at this lol :/

OpenStudy (anonymous):

*see

OpenStudy (aaronq):

haha it's weird, because you wrote the correct value a few sentences earlier, in the same post.

OpenStudy (aaronq):

so for the last question, you need to determines the amount of moles, so: just use \(n=\dfrac{PV}{RT}\) when P<269 kPa

OpenStudy (aaronq):

at the temperature of the canyon.

OpenStudy (anonymous):

hmm im a little lost what to put :/

OpenStudy (anonymous):

PV = nRT A) V = nRT/P to convert deg C to Kelvin add 273 = -4 + 273 = 269 K use R value 0.08206 Latm/Kmol V =1.40 mol x 0.08206 Latm/Kmol x 269 K / (1.25atm) = 24.7 L B) T = PV/nR V = 488/1000 = 0.488 L 1 atm = 760mmHg to covert mmHg to atm divide by 760 755mmHg x (1 atm/760mmHg) = 0.993 atm use R value 0.08206 Latm/Kmol Plug in the equation and solve for T 1 atm = 101 325 pascals = 101.325 kPa 11.15kPa x (1 atm/ 101.325 kPa) = 0.110 atm

OpenStudy (anonymous):

that is what i have but i feel like i am very very wrong and off

OpenStudy (aaronq):

that is incorrect. First use the conditions at the grand canyon to the determine the amount of gas (moles) you need to have so the pressure remains below 269 kPa. then use those moles and the conditions not at the canyon to determine the new pressure your tires should be at.

OpenStudy (anonymous):

okay wait, that is a little confusing to me, I don't know how to get that particular pressure that is "below" I guess that is where i am getting confused

OpenStudy (aaronq):

use 268.9 kPa

OpenStudy (aaronq):

in real life you would use much lower value.

OpenStudy (anonymous):

wait isnt that what i did? I used that number but w/o the .9

OpenStudy (anonymous):

ohhh okay lol i get it now thank yo so so much :)

OpenStudy (aaronq):

nope, you copied n pasted from here http://answers.yahoo.com/question/index?qid=20131101133503AAIZZao but it's not the right question :S

OpenStudy (anonymous):

that answer i got was from my sister :/ i guess she got it form there grrr i am sorry for that she always looks things up if she doesnt get them.

OpenStudy (anonymous):

i guess that would make sense as to why i am confused.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i asked her for help and she seemed to know what she talking about

OpenStudy (aaronq):

it's fine to look things up, but don't take things blindly, check em out for yourself :)

OpenStudy (anonymous):

yeah i know but i get mad when she lookes things up for me b/c i get into trouble for puting them as my answers

OpenStudy (anonymous):

because it can be considered an act of plagiarism in my school.

OpenStudy (anonymous):

but i will definatley take your advice about checking them out for myself before making them my answer

OpenStudy (aaronq):

good stuff. yeah, copying and pasting can get you kicked out of school :S, be careful and good luck !

OpenStudy (anonymous):

okay and thank you so much for your help! :) also do you think you could fan me back?? please so i can pm you when your available you have it set to wear those who are not your fans can not pm you

OpenStudy (aaronq):

oh yeah, sure! sorry, i didn't notice you added me. i get a lot of notifications :S